OpenStudy (aravindg):

A boy standing on a stationary lift (open from above) throws a ball upwards with the maximum initial speed he can, equal to 49 m/s. How much time does the ball take to return to his hands? If the lift starts moving up with a uniform speed of 5 m/s and the boy again throws the ball up with the maximum speed he can, how long does the ball take to return to his hands?

OpenStudy (aravindg):

u see i got first part

OpenStudy (aravindg):

10 s

OpenStudy (aravindg):

but i dont undestand y we get same answr in 2nd case

OpenStudy (jamesj):

The position of the ball at time time is \[ y(t) = 49 t - \frac{1}{2}gt^2 \] \[ = 49 t - 4.9t^2 \] \[ = 0 \] when \( t = 0, 10 \) seconds, right. Now in the second case, the speed of the ball is (49 + 5) m/s = 54 m/s, so it's position is \[ y_{ball}(t) = 54 t - 4.9t^2 \] ... but ...

OpenStudy (jamesj):

the position of the boy is \[ y_{boy}(t) = 5t \] Hence the question is: for what t is \[ y_{ball}(t) = y_{boy}(t) \] Solve that equation and you'll get back \( t = 0, 10 \) seconds.

OpenStudy (aravindg):

hw u got pos of boy as 5t??

OpenStudy (jamesj):

because the lift is moving up at a constant velocity of 5 m/s

OpenStudy (aravindg):

wow i got the old eqn

OpenStudy (aravindg):

:)

OpenStudy (jamesj):

right, exactly

OpenStudy (aravindg):

thx

OpenStudy (aravindg):

so shall i make it general

OpenStudy (aravindg):

that if someone is in a lift and the lift moves with const speed then time to throw and catch the balll is same as when lift is at rest??

OpenStudy (jamesj):

Yes.

OpenStudy (aravindg):

k

OpenStudy (aravindg):

wt if it was not const speed?

OpenStudy (jamesj):

Then you won't have equality.

OpenStudy (jamesj):

Frames of reference moving at a constant speed are called inertial frames. The laws of Newtonian physics work the same in all inertial frames.

OpenStudy (aravindg):

u mean any system with a=0 is an inertial frame?

OpenStudy (jamesj):

Yes

OpenStudy (aravindg):

wow

OpenStudy (aravindg):

wel james hav u seen a type of problem in which a rope is given and we can slide down that and the breaking tension is given and max a with which we can slide down??

OpenStudy (jamesj):

Yes, but I'm going to go now. I'm sure gogind or jemurray or others can help you with that.

OpenStudy (aravindg):

oh bye